In 30 years, Sarah will be 4 times as old as she is now. How old is she now?
Answer by ewatrrr(24785)


Hi
Let x represent Sara's age NOW
Question states*** 30 years from now...
x+30 = 4x
solving for s
30 = 3x
x = 10, Sara's age NOW
CHECKING our Answer***
10 + 30 = 4*10